### v1

# Dark Matter Search Results from 4.2 Tonne-Years of Exposure of the LUX-ZEPLIN (LZ) Experiment J. Aalbers

Dark Matter Search Results from 4.2 Tonne-Years of Exposure
of the LUX-ZEPLIN (LZ) Experiment
J. Aalbers
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
D.S. Akerib
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
A.K. Al Musalhi
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
F. Alder
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
C.S. Amarasinghe
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
A. Ames
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
T.J. Anderson
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
N. Angelides
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
H.M. Araújo
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
J.E. Armstrong
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
M. Arthurs
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
A. Baker
King’s College London, Department of Physics, London WC2R 2LS, UK
S. Balashov
STFC Rutherford Appleton Laboratory (RAL), Didcot, OX11 0QX, UK
J. Bang
Brown University, Department of Physics, Providence, RI 02912-9037, USA
J.W. Bargemann
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
E.E. Barillier
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
University of Zürich, Department of Physics, 8057 Zürich, Switzerland
D. Bauer
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
K. Beattie
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
T. Benson
University of Wisconsin-Madison, Department of Physics, Madison, WI 53706-1390, USA
A. Bhatti
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
A. Biekert
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
T.P. Biesiadzinski
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
H.J. Birch
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
University of Zürich, Department of Physics, 8057 Zürich, Switzerland
E. Bishop
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
G.M. Blockinger
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
B. Boxer
University of California, Davis, Department of Physics, Davis, CA 95616-5270, USA
C.A.J. Brew
STFC Rutherford Appleton Laboratory (RAL), Didcot, OX11 0QX, UK
P. Brás
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
S. Burdin
University of Liverpool, Department of Physics, Liverpool L69 7ZE, UK
M. Buuck
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
M.C. Carmona-Benitez
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
M. Carter
University of Liverpool, Department of Physics, Liverpool L69 7ZE, UK
A. Chawla
Royal Holloway, University of London, Department of Physics, Egham, TW20 0EX, UK
H. Chen
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
J.J. Cherwinka
University of Wisconsin-Madison, Department of Physics, Madison, WI 53706-1390, USA
Y.T. Chin
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
N.I. Chott
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
M.V. Converse
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
R. Coronel
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
A. Cottle
a.cottle@ucl.ac.ukUniversity College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
G. Cox
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
D. Curran
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
C.E. Dahl
Northwestern University, Department of Physics & Astronomy, Evanston, IL 60208-3112, USA
Fermi National Accelerator Laboratory (FNAL), Batavia, IL 60510-5011, USA
I. Darlington
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
S. Dave
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
A. David
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
J. Delgaudio
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
S. Dey
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
L. de Viveiros
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
L. Di Felice
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
C. Ding
Brown University, Department of Physics, Providence, RI 02912-9037, USA
J.E.Y. Dobson
King’s College London, Department of Physics, London WC2R 2LS, UK
E. Druszkiewicz
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
S. Dubey
Brown University, Department of Physics, Providence, RI 02912-9037, USA
S.R. Eriksen
University of Bristol, H.H. Wills Physics Laboratory, Bristol, BS8 1TL, UK
A. Fan
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
S. Fayer
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
N.M. Fearon
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
N. Fieldhouse
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
S. Fiorucci
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
H. Flaecher
University of Bristol, H.H. Wills Physics Laboratory, Bristol, BS8 1TL, UK
E.D. Fraser
University of Liverpool, Department of Physics, Liverpool L69 7ZE, UK
T.M.A. Fruth
The University of Sydney, School of Physics, Physics Road, Camperdown, Sydney, NSW 2006, Australia
R.J. Gaitskell
Brown University, Department of Physics, Providence, RI 02912-9037, USA
A. Geffre
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
J. Genovesi
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
C. Ghag
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
A. Ghosh
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
R. Gibbons
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
S. Gokhale
Brookhaven National Laboratory (BNL), Upton, NY 11973-5000, USA
J. Green
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
M.G.D.van der Grinten
STFC Rutherford Appleton Laboratory (RAL), Didcot, OX11 0QX, UK
J.J. Haiston
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
C.R. Hall
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
T.J. Hall
University of Liverpool, Department of Physics, Liverpool L69 7ZE, UK
S. Han
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
E. Hartigan-O’Connor
Brown University, Department of Physics, Providence, RI 02912-9037, USA
S.J. Haselschwardt
haselsco@umich.eduUniversity of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
M.A. Hernandez
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
University of Zürich, Department of Physics, 8057 Zürich, Switzerland
S.A. Hertel
University of Massachusetts, Department of Physics, Amherst, MA 01003-9337, USA
G. Heuermann
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
G.J. Homenides
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
M. Horn
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
D.Q. Huang
University of California, Los Angeles, Department of Physics & Astronomy, Los Angeles, CA 90095-1547, USA
D. Hunt
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
E. Jacquet
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
R.S. James
Also at The University of Melbourne, School of Physics, Melbourne, VIC 3010, AustraliaUniversity College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
J. Johnson
University of California, Davis, Department of Physics, Davis, CA 95616-5270, USA
A.C. Kaboth
Royal Holloway, University of London, Department of Physics, Egham, TW20 0EX, UK
A.C. Kamaha
University of California, Los Angeles, Department of Physics & Astronomy, Los Angeles, CA 90095-1547, USA
Meghna K.K.
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
D. Khaitan
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
A. Khazov
STFC Rutherford Appleton Laboratory (RAL), Didcot, OX11 0QX, UK
I. Khurana
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
J. Kim
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
Y.D. Kim
IBS Center for Underground Physics (CUP), Yuseong-gu, Daejeon, Korea
J. Kingston
University of California, Davis, Department of Physics, Davis, CA 95616-5270, USA
R. Kirk
Brown University, Department of Physics, Providence, RI 02912-9037, USA
D. Kodroff
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
L. Korley
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
E.V. Korolkova
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
H. Kraus
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
S. Kravitz
University of Texas at Austin, Department of Physics, Austin, TX 78712-1192, USA
L. Kreczko
University of Bristol, H.H. Wills Physics Laboratory, Bristol, BS8 1TL, UK
V.A. Kudryavtsev
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
C. Lawes
King’s College London, Department of Physics, London WC2R 2LS, UK
D.S. Leonard
IBS Center for Underground Physics (CUP), Yuseong-gu, Daejeon, Korea
K.T. Lesko
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
C. Levy
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
J. Lin
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
A. Lindote
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
W.H. Lippincott
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
M.I. Lopes
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
W. Lorenzon
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
C. Lu
Brown University, Department of Physics, Providence, RI 02912-9037, USA
S. Luitz
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085 USA
P.A. Majewski
STFC Rutherford Appleton Laboratory (RAL), Didcot, OX11 0QX, UK
A. Manalaysay
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
R.L. Mannino
Lawrence Livermore National Laboratory (LLNL), Livermore, CA 94550-9698, USA
C. Maupin
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
M.E. McCarthy
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
G. McDowell
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
D.N. McKinsey
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
J. McLaughlin
Northwestern University, Department of Physics & Astronomy, Evanston, IL 60208-3112, USA
J.B. McLaughlin
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
R. McMonigle
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
E. Mizrachi
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
Lawrence Livermore National Laboratory (LLNL), Livermore, CA 94550-9698, USA
A. Monte
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
M.E. Monzani
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
Vatican Observatory, Castel Gandolfo, V-00120, Vatican City State
J.D. Morales Mendoza
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
E. Morrison
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
B.J. Mount
Black Hills State University, School of Natural Sciences, Spearfish, SD 57799-0002, USA
M. Murdy
University of Massachusetts, Department of Physics, Amherst, MA 01003-9337, USA
A.St.J. Murphy
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
A. Naylor
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
H.N. Nelson
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
F. Neves
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
A. Nguyen
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
C.L. O’Brien
University of Texas at Austin, Department of Physics, Austin, TX 78712-1192, USA
I. Olcina
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
K.C. Oliver-Mallory
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
J. Orpwood
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
K.Y Oyulmaz
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
K.J. Palladino
University of Oxford, Department of Physics, Oxford OX1 3RH, UK
J. Palmer
Royal Holloway, University of London, Department of Physics, Egham, TW20 0EX, UK
N.J. Pannifer
University of Bristol, H.H. Wills Physics Laboratory, Bristol, BS8 1TL, UK
N. Parveen
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
S.J. Patton
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
B. Penning
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
University of Zürich, Department of Physics, 8057 Zürich, Switzerland
G. Pereira
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
E. Perry
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
T. Pershing
Lawrence Livermore National Laboratory (LLNL), Livermore, CA 94550-9698, USA
A. Piepke
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
Y. Qie
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
J. Reichenbacher
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
C.A. Rhyne
Brown University, Department of Physics, Providence, RI 02912-9037, USA
A. Richards
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
Q. Riffard
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
G.R.C. Rischbieter
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
University of Zürich, Department of Physics, 8057 Zürich, Switzerland
E. Ritchey
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
H.S. Riyat
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
R. Rosero
Brookhaven National Laboratory (BNL), Upton, NY 11973-5000, USA
T. Rushton
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
D. Rynders
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
D. Santone
Royal Holloway, University of London, Department of Physics, Egham, TW20 0EX, UK
A.B.M.R. Sazzad
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
R.W. Schnee
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
G. Sehr
University of Texas at Austin, Department of Physics, Austin, TX 78712-1192, USA
B. Shafer
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
S. Shaw
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
T. Shutt
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
J.J. Silk
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
C. Silva
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
G. Sinev
South Dakota School of Mines and Technology, Rapid City, SD 57701-3901, USA
J. Siniscalco
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
R. Smith
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
V.N. Solovov
Laboratório de Instrumentação e Física Experimental de Partículas (LIP), University of Coimbra, P-3004 516 Coimbra, Portugal
P. Sorensen
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
J. Soria
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
I. Stancu
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
A. Stevens
University College London (UCL), Department of Physics and Astronomy, London WC1E 6BT, UK
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
K. Stifter
Fermi National Accelerator Laboratory (FNAL), Batavia, IL 60510-5011, USA
B. Suerfu
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
T.J. Sumner
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
M. Szydagis
University at Albany (SUNY), Department of Physics, Albany, NY 12222-0100, USA
D.R. Tiedt
South Dakota Science and Technology Authority (SDSTA), Sanford Underground Research Facility, Lead, SD 57754-1700, USA
M. Timalsina
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
Z. Tong
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
D.R. Tovey
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
J. Tranter
University of Sheffield, Department of Physics and Astronomy, Sheffield S3 7RH, UK
M. Trask
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
M. Tripathi
University of California, Davis, Department of Physics, Davis, CA 95616-5270, USA
A. Usón
University of Edinburgh, SUPA, School of Physics and Astronomy, Edinburgh EH9 3FD, UK
A. Vacheret
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
A.C. Vaitkus
Brown University, Department of Physics, Providence, RI 02912-9037, USA
O. Valentino
Imperial College London, Physics Department, Blackett Laboratory, London SW7 2AZ, UK
V. Velan
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
A. Wang
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
Kavli Institute for Particle Astrophysics and Cosmology, Stanford University, Stanford, CA 94305-4085, USA
J.J. Wang
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
Y. Wang
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
J.R. Watson
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
University of California, Berkeley, Department of Physics, Berkeley, CA 94720-7300, USA
L. Weeldreyer
University of Alabama, Department of Physics & Astronomy, Tuscaloosa, AL 34587-0324, USA
T.J. Whitis
University of California, Santa Barbara, Department of Physics, Santa Barbara, CA 93106-9530, USA
K. Wild
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
M. Williams
University of Michigan, Randall Laboratory of Physics, Ann Arbor, MI 48109-1040, USA
W.J. Wisniewski
SLAC National Accelerator Laboratory, Menlo Park, CA 94025-7015, USA
L. Wolf
Royal Holloway, University of London, Department of Physics, Egham, TW20 0EX, UK
F.L.H. Wolfs
University of Rochester, Department of Physics and Astronomy, Rochester, NY 14627-0171, USA
S. Woodford
University of Liverpool, Department of Physics, Liverpool L69 7ZE, UK
D. Woodward
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
C.J. Wright
University of Bristol, H.H. Wills Physics Laboratory, Bristol, BS8 1TL, UK
Q. Xia
Lawrence Berkeley National Laboratory (LBNL), Berkeley, CA 94720-8099, USA
J. Xu
Lawrence Livermore National Laboratory (LLNL), Livermore, CA 94550-9698, USA
Y. Xu
University of California, Los Angeles, Department of Physics & Astronomy, Los Angeles, CA 90095-1547, USA
M. Yeh
Brookhaven National Laboratory (BNL), Upton, NY 11973-5000, USA
D. Yeum
University of Maryland, Department of Physics, College Park, MD 20742-4111, USA
W. Zha
Pennsylvania State University, Department of Physics, University Park, PA 16802-6300, USA
E.A. Zweig
University of California, Los Angeles, Department of Physics & Astronomy, Los Angeles, CA 90095-1547, USA
Abstract
We report results of a search for nuclear recoils induced by weakly interacting massive particle (WIMP) dark matter using the LUX-ZEPLIN (LZ) two-phase xenon time projection chamber. This analysis uses a total exposure of tonne-years from 280 live days of LZ operation, of which tonne-years and 220 live days are new. A technique to actively tag background electronic recoils from 214Pb decays is featured for the first time. Enhanced electron-ion recombination is observed in two-neutrino double electron capture decays of 124Xe, representing a noteworthy new background. After removal of artificial signal-like events injected into the data set to mitigate analyzer bias, we find no evidence for an excess over expected backgrounds. World-leading constraints are placed on spin-independent (SI) and spin-dependent WIMP-nucleon cross sections for masses 9 GeV/. The strongest SI exclusion set is cm2 at the 90% confidence level at a mass of 36 GeV/, and the best SI median sensitivity achieved is cm2 for a mass of 40 GeV/.
Dark Matter, Direct Detection, Xenon
Astrophysical observations provide strong evidence for the existence of dark matter and its dominance of the matter density of the Universe [1, 2, 3, 4]. Well-motivated as a thermal relic, the weakly interacting massive particle (WIMP) is a leading hypothesis for dark matter, arising in a number of beyond-the-standard-model theories [5, 6, 7]. The LUX-ZEPLIN (LZ) collaboration operates the world’s largest two-phase xenon time projection chamber (TPC), with the primary aim of directly detecting WIMPs through their interactions with xenon nuclei. This technology has produced the most stringent constraints on spin-independent (SI) WIMP-nucleon interactions for 5 GeV/ masses, with all three current-generation experiments having reported first results [8, 9, 10]. In this Letter we present the latest WIMP analysis from LZ, combining a new 220 live-day exposure (WS2024) with the 60 live-day exposure from the first result (WS2022) [8] to perform the most sensitive direct search for WIMPs to date.
The LZ experiment [11, *lztdr, *LZExperiment] is located 4850 ft underground, shielded by a 4300 meter water equivalent rock overburden in the Davis Cavern at the Sanford Underground Research Facility (SURF) in Lead, South Dakota, USA. Its TPC, containing seven tonnes of active liquid xenon (LXe), is surrounded by a dual-detector veto system comprising a two-tonne LXe “Skin,” designed to tag rays, and an outer detector (OD) containing 17 tonnes of gadolinium-loaded liquid scintillator (GdLS), optimized for the detection of neutrons.
Interactions in the LXe target that fills the TPC create light in the form of prompt vacuum ultraviolet (VUV) scintillation (S1) and delayed electroluminescence (S2), both observed by the top and bottom photomultiplier tube (PMT) arrays. The S2 signal is generated by ionization electrons as they traverse a thin layer of gaseous xenon above the liquid, having been drifted upwards from the interaction site under an applied electric field and extracted from the liquid phase. The measured time interval between the S1 and the S2 pulses, referred to as the “drift time”, corresponds to the depth, , of the interaction, while the hit pattern of S2 light in the top PMT array allows for position reconstruction in the transverse plane. The ratio of S2 to S1 signal sizes is used to discriminate between background electronic recoils (ERs) and nuclear recoils (NRs), the latter of which is expected for WIMP-nucleus scatters. Background events are further rejected if they exhibit multiple scatters (MS) rather than a single scatter (SS) in the TPC, or through the presence of coincident signals in the Skin and OD.
The LXe target is contained in a vacuum-insulated cryostat, and is continuously circulated for gas-phase purification to remove electronegative impurities, and to regulate the liquid level. The cooling and circulation systems allow for control of the LXe flow pattern inside the cryostat, which influences the spatial distribution and movement of radioisotopes in the TPC volume. Two distinct flow configurations can be attained, referred to as the “high-mixing” and “low-mixing” states. In the high-mixing state, the liquid flow results in spatially homogeneous distributions of injected calibration sources. In the low-mixing state, there exist distinct regions of laminar flow throughout the detector, and coincident 222RnPo decays can be reliably used to map the LXe velocity field. The ability to model the flow in this state is exploited in the development of a “radon tag” to target the prominent 214Pb ER background, discussed later.
Cathode, gate, and anode electrodes establish the TPC drift field at 97 V/cm and the liquid-gas extraction field at 3.4 kV/cm (at the radial center, below the liquid surface) for the WS2024 data-taking campaign. With respect to WS2022, the extraction field was decreased to reduce the dead time incurred from spurious electron and photon emission, and the cathode voltage was lowered in response to the onset of persistent light emission localized to the Skin region below the cathode.
This analysis incorporates 220.0 live days collected in the new electrode bias configuration between March 27, 2023 and April 1, 2024. Of this WS2024 live time, 40.9 live days are included from when the detector is in the high-mixing state, due to proximity to injected calibrations, with the remaining 179.1 live days taken in the low-mixing state. The xenon gas pressure throughout the run is 1.860 bar(a) with variation at the level of 0.03%. The level of convective mixing is controlled by adjusting the temperature of the liquid returning to the bottom of the TPC from 177.9 K (high-mixing state) to 174.2 K (low-mixing state), with variations at the level of 0.02% for each. Changing from high-mixing to low-mixing state increases the maximum drift time from 1045 s to 1050 s and reduces the overall liquid height by about 0.2 mm. The consequent 1% increase in S2 signal size is corrected for in analysis.
Data acquisition uses a trigger with a digital filter designed primarily to capture S2 pulses. The trigger was optimized for WS2024 to increase the sensitivity to low-energy events of interest to the WIMP analysis, achieving an efficiency of 95% for S2 pulses corresponding to 3.5 extracted electrons. Waveforms from the TPC, Skin, and OD PMTs are recorded with a window extending from 2.2 ms before the trigger time to 2.5 ms after it, sampled at 100 MHz. Further details of the data acquisition and signal chain are described in Ref. [14].
Event waveforms are processed to find and categorize pulses and interactions. Pulse sizes are normalized by the time-dependent size of single VUV photon pulses measured in each channel, and are reported as the number of photons detected (phd) [15, 16]. Events are categorized as SS if they have a single S1 preceding a single S2. The SS identification efficiency is dependent on the S1 and S2 pulse classification efficiencies. Ionization electrons diffuse as they drift, which can lead to fragmentation of the S2 signal into multiple pulses. The splitting probability is 1% on average across the TPC for S2 pulses relevant to this analysis, as estimated with waveform simulations [17]. For a pulse to be classified as S1, coincident photons must be recorded in at least three PMTs; however, undersized PMT responses and differences in photon transit times can lead to missed S1 pulses. The SS S1 identification efficiency is found to be 90% for all S1 pulse sizes in this analysis, rising to 100% for S1s 35 phd (8 keV). This efficiency is assessed by comparing the event classification of a sample of tritium and deuterium-deuterium (DD) neutron calibration events determined by the event reconstruction algorithms to the corresponding classification obtained from a manual, visual categorization campaign.
Tritium and DD neutrons are among several sources used to calibrate the TPC [18]. Tritium is delivered through radiolabeled methane along with 14C, in an activity ratio of approximately 8:1. The tritiated methane, primarily used to calibrate the ER response, was injected into the TPC while in the high-mixing state to ensure its efficient dispersal and complete removal. Mono-energetic 2.45 MeV neutrons from a collimated DD generator, directed roughly 10 cm below the LXe surface, supply the counterpart for calibrating the NR response. AmLi neutron sources, deployed in source tubes at three different azimuthal positions and three different heights around the TPC, provide additional NR calibration events. Injections of 131mXe and 83mKr at roughly monthly and trimonthly cadences, respectively, in addition to background radon progeny, provide mono-energetic peaks throughout WS2024, used to monitor the stability of the detector response and develop corrections for the normalization of S1 and S2 signals.
S1 and S2 signals exhibit spatial and temporal variations due to factors such as changing detector conditions or non-uniform light collection and electron extraction efficiencies. Corrections are applied to produce standardized values, S1 and S2. An map normalizes the S1 signals to the center of the TPC volume, with correction factors averaging 8% in absolute value. S2 signals are normalized to the radial center of the liquid surface in the plane, in addition to being corrected for the depth-dependent probability of electron attachment to electronegative impurities in the LXe; the average values of these corrections are approximately 13% and 7%, respectively. The responses are also normalized in time, using the tritium calibration period as the reference.
Calibrations also inform the tuning of the nest 2.4.0 [19] model for detector response, in which the scintillation photon and ionization electron gains, and , are key parameters. A scan over viable values for and , performed using tritium data in space and constrained by the energy reconstruction of 131mXe and 83mKr peaks, yields phd/photon and phd/electron. A further tuning to more accurately capture the “ER band,” i.e. the shape of the observed distribution of tritium data, enables nest to reproduce the band median and width to better than 0.2% and 5% in , respectively. The DD data define an analogous “NR band,” for which nest is separately tuned to a similar level of precision via small adjustments in the NR yield parameters. The DD neutron interactions are confined to the top of the detector, where the average electric field is 7% higher than the center. Because the charge yield of 222Rn decays used for spatial corrections is more field-sensitive than NRs, corrections in the DD neutron region overestimate field effects on recombination, leading to a systematic shift in the NR band position as assumed for the full TPC. An upwards shift of 2.4% in S2 is applied to the NR model to compensate for this effect, calculated from electrostatic simulations and nest-based models.
Candidate events for the WIMP analysis are selected if they are classified as SS; they are located within the fiducial volume (FV); they pass veto coincidence, S1 and S2-based, and live-time exclusion cuts, all of which are described below; and they lie within the WIMP region of interest (ROI). The ROI is defined such that the corrected S1 and S2 pulse sizes are in the range of phd and . A lower bound on uncorrected S2 pulse size of 645 phd, equivalent to 14.5 extracted electrons, is also applied. This S2 threshold renders the analysis largely insensitive to 8B coherent elastic neutrino-nucleus scattering (CENS), a search for which will be reported in a future publication.
Figure 1: Data from WS2024, in black with all analysis cuts applied, and in gray outside the FV without veto coincidence cuts. Red crosses and blue circles represent events that are vetoed by the prompt and delayed veto coincidence cuts, respectively. The dashed line shows the active volume, averaged over azimuth, and the two solid lines depict the FV, at the azimuths of its smallest and largest radial extent. The drift field affects the position reconstruction such that the observed radius of the TPC varies as a function of both drift time and azimuthal angle.
A FV, illustrated in Fig. 1, is defined to suppress external backgrounds. Drift time limits are placed at 71 s and 1034 (1030) s in the low(high)-mixing state, corresponding to distances of 9.9 cm below the gate and 2.2 cm above the cathode, respectively. Events within 6.0 cm of the TPC field cage resistors [13] in are also excluded, as these resistors have elevated radioactivity. The radial fiducial boundary is defined such that backgrounds originating from the TPC wall, which can be misreconstructed radially inwards due to finite position resolution, have a negligible expectation of events for WS2024. This radial boundary follows a contour of measured wall background event rate, and is a function of both depth and azimuthal angle. The fiducial mass, determined using the fraction of tritium events in the full active volume found within the FV, is tonnes.
Veto coincidence cuts reject events that are accompanied by activity in the Skin and/or OD, and are divided into prompt and delayed categories (see Fig. 1). Prompt cuts target rays and fast neutron scatters, removing events with signals in the Skin (OD) of size phd ( phd) within 0.25 s (0.3 s) of the S1. Delayed veto cuts target neutrons, which thermalize then capture, typically on Gd in the OD with the release of an 8–9 MeV -ray cascade. These cuts reject events with pulses in the OD (Skin) of energy keV ( keV) that occur within a 600 s time window after the S1. The delayed veto cuts remove 3% of the collected live time due to coincidences between unrelated interactions in the TPC and the veto detectors. The veto efficiency for AmLi calibration neutrons, combining prompt and delayed cuts, is measured to be %. Radiogenic neutrons can be vetoed more effectively as they are typically higher in energy than those from AmLi, and are often accompanied by other neutrons and rays. Simulations tuned on the AmLi calibration data indicate that the efficiency to reject background neutrons is %, with the uncertainty driven by residual discrepancies between the simulations and AmLi data.
Cuts based on S1 and S2 pulse characteristics target accidental coincidence events, in which the random pairing of pulses classified as S1s and S2s mimic true SS interactions. Several sources give rise to isolated S1- and S2-classified pulses, such as interactions in light- or charge-insensitive regions of the detector, spurious Cherenkov photon emissions, or delayed emissions of electrons [20, 21]. The cuts primarily exploit the shapes of these pulses in assessing the validity of the S1-S2 pair. The rejection power of the cuts is evaluated on a sideband of known accidental coincidence events with unphysical drift times (1100 s) and is 99.5% for all such events in the WIMP ROI.
Live time exclusions discount periods with higher-than-average incidences of accidental coincidences due to elevated pulse rates, which generally occur in the wake of high-energy interactions. A time hold-off is imposed after large S2 signals, dependent on their size, to avoid subsequent delayed ionization or correlated light emission. The hold-off calculation is optimized with respect to that used in WS2022 so that it cumulatively removes 10% of the recorded live time in WS2024, as opposed to 30% in WS2022. Accounting for all live-time losses and exclusions, 81.5% of the collected live time of WS2024 remains for analysis, totaling 220.0 live days.
The cumulative effect of the data selection, event detection, and reconstruction efficiencies on the NR signal efficiency can be seen in Fig. 2, where the cut acceptances are assessed on a combination of tritium and neutron calibration data sets. Uncertainties arise from the combined statistical error on the SS efficiency and data analysis cuts, as well as a systematic error equal to the variation in efficiency of data analysis cuts as evaluated on the different calibration data sets.
Figure 2: Energy-dependent NR signal efficiency in WS2024 after the application, in sequence, of the S2 trigger (blue); SS reconstruction and analysis cuts (orange); S2 ROI (green) followed by the S1 ROI (black). The inset shows the low-energy behavior, with the dotted line at 5.4 keV marking 50% efficiency. The uncertainty band (gray) is assessed using AmLi and tritium data as discussed in the text.
To mitigate analyzer bias, artificial, WIMP-like events are injected at random times into the data pipeline, a process referred to as “salting.” Salt events are pre-generated by combining S1 and S2 pulses from sequestered tritium and AmLi calibration data sets. The pulse sizes for a given event are chosen to best match those expected from an NR energy deposition drawn from a parent distribution consisting of the sum of an exponential (WIMP-like) and a flat spectrum. The salt injection rate and parameters of the parent energy distribution were generated randomly within predefined bounds and kept hidden for the entirety of the analysis. Eight salt events were injected, of which one was removed by analysis cuts, consistent with the NR signal efficiency. The other seven events were revealed as salt following the final definition of the data selection criteria and the likelihood models.
The 1220 events which remain in the 3.3 tonne-year exposure of WS2024 after the application of all data selections and salt removal are shown in Figs. 1 and 3. These events are analyzed using a likelihood model in the observables. Signal and background model distributions are generated using nest and event simulations [22], with the exception of the accidental coincidence background model, which follows a data-driven method identical to that used for the WS2022 result [23]. The primary backgrounds and their treatment are similar to those of the WS2022 analysis, with the expected counts for each background source given in Table 1. Notable differences are now described.
Two sources of background NR events are considered: those originating from neutrino interactions and those from radioactive decays in detector components. The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos. Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and OD detectors. Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data. As discussed later, the sideband is incorporated directly into the final statistical analysis; however, a sideband-only fit returns an expectation of SS counts, as shown in Table 1. Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
Electron capture (EC) and double electron capture (DEC) decays deposit their energy via X ray and Auger cascades in a more localized profile than a particle of the equivalent energy, leading to higher ionization densities and enhanced electron-ion recombination. As a result, EC and DEC events look more NR-like than decays, as reported in Ref. [24]. Preliminary measurements of this effect for ECs in LZ were performed with sidebands of Skin-tagged and multiple scatter decays of 127Xe and 125Xe, isotopes that have been produced in situ via neutron activation. These yield a charge-yield ratio of for 5.2 keV L-shell EC decays relative to decays of the same energy.
Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [25], consistent with that in Ref. [26]. The expected number of 124Xe counts in the ROI is estimated assuming an abundance of [27] and branching ratios for the LL (10.00 keV) and LM (5.98 keV) decay modes of 1.4% and 0.8%, respectively. The LM charge suppression is modeled with the same ratio of 0.87 as used for single EC decays. The overlap of two L-shell cascades in LL-shell capture should result in additional recombination. As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [28, 29] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field. The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the expectation for a doubled ionization density from the TI box model (0.65).
The largest ER background contribution in this exposure comes from -emitting isotopes in the LXe bulk. Xenon sampling measurements indicate a concentration of ppb Ar g/g and an exposure-averaged concentration of ppq Kr g/g. These are used to estimate the contributions of 39Ar and 85Kr, respectively. Exposure in the high-mixing state includes residual tritium and 14C activity following calibrations, and therefore both their beta spectra are incorporated into the background model for this period.
Radon emanates from materials exposed to xenon, dispersing in the active volume. With a measured FV activity of Bq/kg, 214Pb decays from the 222Rn chain, in particular those to the ground state, form the dominant ER background. A method to identify events with high likelihood of being a 214Pb decay — the radon tag — is developed using simulations of the movement of daughter atoms and ions built from observed flow vectors between 222Rn-218Po -decay pairs. A tag volume in which it is probable to find a 214Pb decay is defined for each identified 218Po alpha decay, co-moving with the projected flow streamline for 81 min or approximately 3 the 214Pb half-life. The radon tag is applied in the low-mixing state, where the flow can be modelled, and when data are available for a continuous 81 min period such that tag volumes can be constructed. For WS2024, the radon tag identifies 15% of the tag-active exposure as 214Pb-rich, containing on average % of the 214Pb, estimated directly by measurements of its excited-state decays. The effective 214Pb activity in the untagged remainder of the tag-active exposure is thus Bq/kg. Both the radon-tagged and untagged data sets are used in the final WIMP analysis; the former can be seen alongside all the WIMP search data in the bottom panel of Fig. 3. The separation into 214Pb-rich and 214Pb-poor exposures offers the benefit of being able to better identify and constrain other background contributions, 124Xe in particular.
Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom). 214Pb is at 60% and disperse backgrounds at 15% of their full FV activity in the radon-tagged data set; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band. Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively. Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively. Gray lines show contours of constant ER-equivalent (keVee) and NR-equivalent (keVnr) energy.
Table 1: The expected and best-fit counts for different sources in the 3.3 tonne-year exposure of WS2024, including a 40 GeV/ WIMP signal. Flat -decay components have been separated based on whether they are affected by the flow state or the radon tag, and the neutron counts are derived in situ. “Det. s” refers to -ray contributions from detector materials, whose Compton spectra are also flat in the ROI.
Source
Pre-fit Expectation
Fit Result
[…]
743
880
733
34 0
85Kr + 39Ar s + det. s
162
220
161
21 0
Solar ER
102
6 0
102
6 0
212Pb + 218Po s
62.7
7.5 0
63.7
7.4 0
Tritium+ 14C s
58.3
3.3 0
59.7
3.3 0
[…]
55.6
8.3 0
55.8
8.2 0
[…]
19.4
3.9
21.4
3.6
[…]
3.2
0.6
2.7
0.6
Accidental coincidences
2.8
0.6
2.6
0.6
NR
0.12
0.02
0.12
0.02
8B NR
0.06
0.01
0.06
0.01
Detector neutrons
00111The expected number of neutron events results from a fit to the sample of veto detector-tagged events. This expectation is not explicitly used in the final combined fit as this sample is included directly in the likelihood, as described in the text.
40 GeV/ WIMP
Total
1210
91
1203
42 0
The data are divided into six mutually exclusive samples, which are simultaneously fit using an unbinned, extended maximum likelihood. Five of these apply to the new 220 live-day exposure and describe: (1) events acquired in the high-mixing state; (2) events acquired in the low-mixing state when the radon tag was inactive; events acquired in the low-mixing state that are (3) radon tagged, and (4) radon untagged; and (5) events tagged by the Skin or OD. The nuisance parameters, which float within their uncertainties, are the expected number of events for each background source, listed in Table 1; the radon tagging efficiency; the neutron veto tagging efficiency; and the
124Xe , bounded between 0.65 and 0.87. Parameters characterizing 222Rn and 220Rn daughter backgrounds, tritium and 14C are independent in the high- and low-mixing state exposures. The sixth sample in the likelihood represents WS2022, containing the data and models from the first, 60 live-day WIMP search result [8].
Figure 4: Comparison between data and the best-fit model for WS2024 corresponding to the values shown in Table 1 for a background plus a 40 GeV/ WIMP signal fit in a projection along the ER band median, normalized to the band width. The 40 GeV/ WIMP component is shown with an amplitude equal to its power-constrained 90% C.L. upper limit. The upper and lower panels select the S1 ranges 3–40 phd and 40–80 phd with values of 0.91 and 0.27, respectively. Shaded bands depict the central interval containing 68% of the combined systematic and statistical uncertainties of the model.
The best-fit number of WIMPs at all masses tested (between 9 GeV/ and 100 TeV/) for the combined WS2024+WS2022 analysis is zero. The goodness of fit of the background-only model is assessed in binned 1D projections of the data and model into S1c, log10(S2c), reconstructed energy, distance from the ER band median, as well as the 2D space of . All samples show excellent model-data agreement, passing a Holm-Bonferroni test [30] with a significance level of 0.05. Figure 4 shows the WS2024 data and model in the 1D projection along the ER band median, illustrating the level of agreement through the tails of the distribution with the inclusion of the modified 124Xe response. The fitted values of the radon tagging efficiency and the 124Xe are consistent across all masses, at % and , respectively, in agreement with the pre-fit expectations.
Figure 5 shows the 90% confidence level upper limit on the spin-independent WIMP-nucleon cross section as a function of mass following inference performed with a two-sided, unbinned profile likelihood ratio test statistic [31]. The limit is power constrained at all WIMP masses to below the median expectation following the recommendation of Ref. [*[][;Hereweemployapowerconstraintwhichusestheexpectedsensitivityoftheexperimentassumingnosignalispresent.]baxter2021recommended]. The fluctuation, and therefore the size of the constraint, is largest between 20 and 50 GeV/. This is a consequence of the inclusion of the WS2022 data from Ref. [8], which exhibit an underfluctuation in this region. The limit for the WS2024 data alone can be found in the Supplemental Material, with an approximate fluctuation across the entire mass range, primarily attributed to an underfluctuation of the accidental coincidence background in the region of largest overlap with expected WIMP distributions. The minimum of the limit curve is cm2 at 36 GeV/ while that of the median expected limit is cm2 at 40 GeV/c2. The best median observation potential is cm2 at 40 GeV/c2. Results of searching for evidence of spin-dependent WIMP-nucleon couplings are discussed in the Appendix.
Figure 5:
Upper limits (90% C.L.) on the spin-independent WIMP-nucleon cross section as a function of WIMP mass from the combined WS2024+WS2022 analysis (280.0 live days) are shown with a solid black line, with a power constraint applied. The gray dot-dash line shows the limits without the power constraint; green and yellow regions show the range of expected upper limits from 68% and 95% of background-only experiments, while the dashed black line indicates the median expectation, obtained with post-fit background estimates. The median observation potential from the post-fit model is shown as a dotted black line. Also shown are limits from WS2022 only [8], PandaX-4T [10], LUX [33], all power constrained to ; XENONnT [9], reinterpreted with a power constraint; XENON1T [34], and DEAP-3600 [35].
In summary, LZ has achieved limits on SI WIMP-nucleon interactions that surpass previous best exclusions by a factor of four or more for WIMP masses 9 GeV/. A radon tag to target the dominant ER background from 214Pb decays has been demonstrated, with potential for optimization that will allow it to have greater impact on the sensitivities of future searches, particularly for physics with potential ER signals. The enhanced electron-ion recombination of 124Xe LL-shell DEC events was noted for the first time, and will be further investigated in future analyses. The experiment continues to take salted data towards a target 1000-day live time that will enable more sensitive searches for WIMP interactions and other new phenomena.
Acknowledgements - The research supporting this work took place in part at the Sanford Underground Research Facility (SURF) in Lead, South Dakota. Funding for this work is supported by the U.S. Department of Energy, Office of Science, Office of High Energy Physics under Contract Numbers DE-AC02-05CH11231, DE-SC0020216, DE-SC0012704, DE-SC0010010, DE-AC02-07CH11359, DE-SC0015910, DE-SC0014223, DE-SC0010813, DE-SC0009999, DE-NA0003180, DE-SC0011702, DE-SC0010072, DE-SC0006605, DE-SC0008475, DE-SC0019193, DE-FG02-10ER46709, UW PRJ82AJ, DE-SC0013542, DE-AC02-76SF00515, DE-SC0018982, DE-SC0019066, DE-SC0015535, DE-SC0019319, DE-SC0024225, DE-SC0024114, DE-AC52-07NA27344, & DE-SC0012447. This research was also supported by U.S. National Science Foundation (NSF); the UKRI’s Science & Technology Facilities Council under award numbers ST/W000490/1, ST/W000482/1, ST/W000636/1, ST/W000466/1, ST/W000628/1, ST/W000555/1, ST/W000547/1, ST/W00058X/1, ST/X508263/1, ST/V506862/1, ST/X508561/1, ST/V507040/1, ST/W507787/1, ST/R003181/1, ST/R003181/2, ST/W507957/1, ST/X005984/1, ST/X006050/1; Portuguese Foundation for Science and Technology (FCT) under award numbers PTDC/FIS-PAR/2831/2020; the Institute for Basic Science, Korea (budget number IBS-R016-D1); the Swiss National Science Foundation (SNSF) under award number 10001549. This research was supported by the Australian Government through the Australian Research Council Centre of Excellence for Dark Matter Particle Physics under award number CE200100008. We acknowledge additional support from the UK Science & Technology Facilities Council (STFC) for PhD studentships and the STFC Boulby Underground Laboratory in the U.K., the GridPP [36, 37] and IRIS Collaborations, in particular at Imperial College London and additional support by the University College London (UCL) Cosmoparticle Initiative, and the University of Zurich. We acknowledge additional support from the Center for the Fundamental Physics of the Universe, Brown University. K.T. Lesko acknowledges the support of Brasenose College and Oxford University. The LZ Collaboration acknowledges the key contributions of Dr. Sidney Cahn, Yale University, in the production of calibration sources. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231. We gratefully acknowledge support from GitLab through its GitLab for Education Program. The University of Edinburgh is a charitable body, registered in Scotland, with the registration number SC005336. The assistance of SURF and its personnel in providing physical access and general logistical and technical support is acknowledged. We acknowledge the South Dakota Governor’s office, the South Dakota Community Foundation, the South Dakota State University Foundation, and the University of South Dakota Foundation for use of xenon. We also acknowledge the University of Alabama for providing xenon. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) license to any Author Accepted Manuscript version arising from this submission. Finally, we respectfully acknowledge that we are on the traditional land of Indigenous American peoples and honor their rich cultural heritage and enduring contributions. Their deep connection to this land and their resilience and wisdom continue to inspire and enrich our community. We commit to learning from and supporting their effort as original stewards of this land and to preserve their cultures and rights for a more inclusive and sustainable future.
Appendix A Appendix - Spin Dependent Results
[…]
Figure 6: The 90% confidence-level upper limit (solid-black curve) using the mean of the nuclear structure functions from [39], allowing for like-to-like comparison with other results.
The gray band indicates the uncertainty due to nuclear modeling across models from [39, 40, 41] and applies similarly to the limit from all xenon-based experiments. The dotted line shows the median observation significance from the post-fit model.
Also shown are the WS2022 only [8], XENONnT [9] (reinterpreted with a power constraint), XENON1T [42], PandaX-II [43], and LUX [44] limits.Figure 7: The 90% confidence-level upper limit (solid-black curve) using the mean of the nuclear structure functions from [39], allowing for like-to-like comparison with other results.
The gray band indicates the uncertainty due to nuclear modeling across models from [39, 40, 41] and applies similarly to the limit from all xenon-based experiments.
The dotted line shows the median observation significance from the post-fit model.
Also shown are the PICO-60 [45], LZ WS2022 only [8], XENONnT [9] (reinterpreted with a power constraint), XENON1T [42], PandaX-II [43], and LUX [44] limits. The PICO-60 result relies on WIMP-scattering on the spin of the unpaired proton of 19F with minimal uncertainty.
The best-fit number of WIMP events is zero at all tested masses, from 9 GeV/ to 100 TeV/. The black line of Figure 6 shows the 90% confidence level nominal upper limit on the WIMP-neutron spin-dependent cross section as a function of mass, while the gray band shows the nuclear structure function uncertainty on the limit. The minimum of the limit curve occurs at 46 GeV/ at a cross section of (median expected limit at at 40 GeV/), and a power constraint is applied at all tested masses. Figure 7 shows the same for the WIMP-proton spin-dependent cross section. The minimum of the limit curve is at 40 GeV/ at a cross section of , and a power constraint is applied between 9 GeV/ and 91 GeV/. The minimum of the median expected limit on is and occurs at 40 GeV/.
Faham et al. [2015]C. Faham, V. Gehman, A. Currie, A. Dobi, P. Sorensen, and R. Gaitskell, J. Instrum. 10 (09), P09010.
López Paredes et al. [2018] B. López Paredes, H. Araújo, F. Froborg, N. Marangou, I. Olcina, T. Sumner, R. Taylor, A. Tomás, and A. Vacheret, Astropart. Phys. 102, 56 (2018).
Temples et al. [2021]D. J. Temples, J. McLaughlin, J. Bargemann, D. Baxter, A. Cottle, C. Dahl, W. Lippincott, A. Monte, and J. Phelan, Phys. Rev. D 104, 112001 (2021).
de Laeter et al. [2003a]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure and Applied Chemistry 75, 683 (2003a).
Thomas and Imel [1987]J. Thomas and D. Imel, Physical Review A 36, 614 (1987).
Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv preprint arXiv:2211.10726 (2022).
de Laeter et al. [2003b]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure Appl. Chem. 75, 683 (2003b).
Figure S1 shows the WS2024 and WS2022 events passing all data selection criteria grouped in their respective sample in the likelihood. Table S1 lists the exposure for each sample in tonne-years.
Figure S1:
Panels show the events used in the WS2024+WS2022 combined analysis organized into the various samples described in the main text. Model contours are the same as those in Fig. 3. Red contours in the veto tagged panel display the regions containing 68% and 95% of the distribution of tagged neutron backgrounds.
Table S1: Exposures of the six samples used in the likelihood given in tonne-years.
High-Mixing
Radon Tag
Radon
Radon
Veto
WS2022
State
Inactive
Tagged
Untagged
Tagged
Only
0.6
0.6
0.3
1.8
n/a
0.9
Appendix C Details of the Detector Response Model
The LZ detector and xenon response models are implemented in a nest-based application that includes effects such as curved electron drift paths from field non-uniformities, finite position reconstruction resolution in the transverse and longitudinal directions, and position-dependence of the S1 and S2 signal sizes. The key detector parameters of the model are provided in Table S2.
Table S2: Relevant nest detector parameters. The parameters in the top-half of the table are input parameters, while the bottom-half are results of nest calculations.
Parameter
WS2024
WS2022
[phd/photon]
0.112
0.114
[phd/photon]
0.076
0.092
Effective GXe Extraction Field [kV/cm]
8.3
8.4
Single Electron Size [phd]
46.9
58.5
Extraction Efficiency [%]
72.5
80.5
[phd/electron]
34.0
47.1
A header file for nest v2.4.0 that will reproduce the ER and NR response models used in this analysis is available online.
Note that NEST has multiple models for extraction efficiency versus gaseous xenon (GXe) electric field strength due to conflicting measurements in the literature [46], and the default model provides the lower bound of the extraction efficiency for a given field strength. In order to reproduce the observed single electron size and measured extraction efficiency in LZ with this default model, a higher effective value for the GXe field magnitude is used, and thus the value reported in Table S2 does not represent a physical field inside the LZ TPC. Extraction efficiency is measured as the ratio between the best-fit electroluminescence gain () value and the observed single electron signal size.
In addition to the parameters below, the ER and NR mean yield and fluctuations models have been tuned to minimize the discrepancy between NEST and LZ calibration data. The ER mean yield parameters are provided in Table S3, and the NR parameters are shown in Table S4. Nomenclature for the individual parameters follows from [29]. There are detailed instructions for implementing these changes in the provided file referenced above.
Table S3: Default and best-fit model parameters for the nest ER mean charge yield model. The LZ values were found by fitting to the 3H and 14C band means. See Eq. 6 of [29] for more information regarding the individual .
Parameter
nest Default at 97 V/cm
LZ Value
[keV-1]
14.6
12.5
[keV-1]
77.3
85.0
[keV]
0.8
0.6
2.1
2.1
[keV-1]
19.3
25.7
[keV-1]
0
0
[keV]
78.0
59.7
4.3
3.7
0.33
0.28
0.08
0.11
Table S4: Model parameters for the nest NR mean light and charge yields model. The LZ values were found by fitting to the calibrated DD band mean. See Eqs. (10), (12), and (13) of [29] for more information regarding the individual parameters. The last two parameters ( and ) are not described in [29], but are available as free parameters in nest, controlling the asymmetry of the sigmoidal energy dependence for the charge and light yields, respectively. See the NEST function GetYieldNR(…) for the usage of and in calculating NR yields [19].
Parameter
nest Default
LZ Value
[keV1/β]
11.0
10.2
1.1
1.1
[(V/cm)1/δ]
0.048
0.050
-0.0533
-0.0533
[keV]
12.6
12.5
[keV]
0.30
0.29
2.0
1.9
[keV]
0.30
0.32
2.0
2.1
0.50
0.51
1.0
0.996
1.0
0.999
Appendix D WS2024-Only Limit
Figure S2 shows the SI WIMP-nucleon cross section upper limit obtained using the data and models pertaining to the WS2024, 3.3 tonne year exposure only.
Figure S2: Upper limits (90% C.L.) on the spin-independent WIMP-nucleon cross-section as a function of WIMP mass from WS2024 (220.0 live days) only are shown with a solid black line, with a power constraint applied. The gray dot-dash line shows the limits without the power constraint; green and yellow regions show the range of expected upper limits from 68% and 95% of background-only experiments, while the dashed black line indicates the median expectation, obtained with post-fit background estimates. The median observation significance from the post-fit model is shown as a dotted black line. Also shown are limits from WS2022 only [8], PandaX-4T [10], LUX [33], all power constrained to ; XENONnT [9], reinterpreted with a power constraint; XENON1T [34], and DEAP-3600 [35].
Appendix E Limits in Terms of Number of WIMP Events
Figure S3 presents the WS2022+WS2024 combined and WS2024-only upper limits in number of events.
Figure S3: The confidence limit (solid line) for the number of WIMP events versus WIMP mass in the spin-independent case for the combined analysis (left) and WS2024 only (right). The dot-dash line shows the limit before applying the power constraint described in the text. The dashed line shows the median sensitivity and the dotted line shows the median discovery sensitivity. The green and yellow bands are the and sensitivity bands.

### v2

@@
 The ratio of S2 to S1 signal sizes is used to discriminate between background electronic recoils (ERs) and nuclear recoils (NRs), the latter of which is expected for WIMP-nucleus scatters.
-Background events are further rejected if they exhibit multiple scatters (MS) rather than a single scatter (SS) in the TPC, or through the presence of coincident signals in the Skin and OD.
+Background events are further rejected if they exhibit multiple scatters (MS) rather than a single scatter (SS) in the TPC, or through the presence of coincident signals in the Skin and/or OD.
 The LXe target is contained in a vacuum-insulated cryostat, and is continuously circulated for gas-phase purification to remove electronegative impurities, and to regulate the liquid level.
@@
 In the low-mixing state, there exist distinct regions of laminar flow throughout the detector, and coincident 222RnPo decays can be reliably used to map the LXe velocity field.
-The ability to model the flow in this state is exploited in the development of a “radon tag” to target the prominent 214Pb ER background, discussed later.
+The ability to model the flow in this state is exploited in the development of a “radon tag” to target the prominent 214Pb ER background (similar to that in Ref.
+[14]), discussed later.
 Cathode, gate, and anode electrodes establish the TPC drift field at 97 V/cm and the liquid-gas extraction field at 3.4 kV/cm (at the radial center, below the liquid surface) for the WS2024 data-taking campaign.
@@
 The neutrino interactions are from CENS of solar (8B and hep) and cosmic-ray-generated atmospheric neutrinos.
-Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and OD detectors.
+Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and/or OD.
 Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data.
@@
 As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [29, 30] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field.
-The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the expectation for a doubled ionization density from the TI box model (0.65).
+The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
 The largest ER background contribution in this exposure comes from -emitting isotopes in the LXe bulk.
@@
 Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom).
-214Pb is at 60% and disperse backgrounds at 15% of their full FV activity in the radon-tagged data set; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band.
+The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band.
 Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively.
@@
 Thomas and Imel [1987]J. Thomas and D. Imel, Physical Review A 36, 614 (1987).
-Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv preprint arXiv:2211.10726 (2022).
+Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv:2211.10726 (2022).
 de Laeter et al. [2003b]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure Appl. Chem. 75, 683 (2003b).
@@
 Appendix C Details of the Detector Response Model
-The LZ detector and xenon response models are implemented in a nest-based application that includes effects such as curved electron drift paths from field non-uniformities, finite position reconstruction resolution in the transverse and longitudinal directions, and position-dependence of the S1 and S2 signal sizes.
-The key detector parameters of the model are provided in Table S2.
+[…]
 Table S2: Relevant nest detector parameters. The parameters in the top-half of the table are input parameters, while the bottom-half are results of nest calculations.
@@
 47.1
-A header file for nest v2.4.0 that will reproduce the ER and NR response models used in this analysis is available online.
 Note that NEST has multiple models for extraction efficiency versus gaseous xenon (GXe) electric field strength due to conflicting measurements in the literature [47], and the default model provides the lower bound of the extraction efficiency for a given field strength.

### v3

@@
 World-leading constraints are placed on spin-independent (SI) and spin-dependent WIMP-nucleon cross sections for masses 9 GeV/.
-The strongest SI exclusion set is cm2 at the 90% confidence level at a mass of 36 GeV/, and the best SI median sensitivity achieved is cm2 for a mass of 40 GeV/.
+The strongest SI exclusion set is cm2 at the 90% confidence level and the best SI median sensitivity achieved is cm2, both for a mass of 40 GeV/.
 Dark Matter, Direct Detection, Xenon
@@
 The LUX-ZEPLIN (LZ) collaboration operates the world’s largest two-phase xenon time projection chamber (TPC), with the primary aim of directly detecting WIMPs through their interactions with xenon nuclei.
-This technology has produced the most stringent constraints on spin-independent (SI) WIMP-nucleon interactions for 5 GeV/ masses, with all three current-generation experiments having reported first results [8, 9, 10].
+This technology has produced the most stringent constraints on spin-independent (SI) WIMP-nucleon interactions for 5 GeV/ masses, with all three current-generation experiments having reported results [8, 9, 10].
 In this Letter we present the latest WIMP analysis from LZ, combining a new 220 live-day exposure (WS2024) with the 60 live-day exposure from the first result (WS2022) [8] to perform the most sensitive direct search for WIMPs to date.
@@
 The consequent 1% increase in S2 signal size is corrected for in analysis.
+Stable electron lifetimes
+averaging 9.3 ms and 13.8 ms were maintained in the
+high-mixing and low-mixing states, respectively.
 Data acquisition uses a trigger with a digital filter designed primarily to capture S2 pulses.
@@
 The SS S1 identification efficiency is found to be 90% for all S1 pulse sizes in this analysis, rising to 100% for S1s 35 phd (8 keV).
-This efficiency is assessed by comparing the event classification of a sample of tritium and deuterium-deuterium (DD) neutron calibration events determined by the event reconstruction algorithms to the corresponding classification obtained from a manual, visual categorization campaign.
+This efficiency is assessed by comparing the event classification of a sample of tritium beta decays and deuterium-deuterium (DD) neutron calibration events determined by the event reconstruction algorithms to the corresponding classification obtained from a manual, visual categorization campaign.
 Tritium and DD neutrons are among several sources used to calibrate the TPC [19].
 Tritium is delivered through radiolabeled methane along with 14C, in an activity ratio of approximately 8:1.
-The tritiated methane, primarily used to calibrate the ER response, was injected into the TPC while in the high-mixing state to ensure its efficient dispersal and complete removal.
+The tritiated methane, primarily used to calibrate the ER response, was injected into the TPC while in the high-mixing state to ensure its efficient dispersal and complete removal by the purification system, with a half-life of 53 hours.
 Mono-energetic 2.45 MeV neutrons from a collimated DD generator, directed roughly 10 cm below the LXe surface, supply the counterpart for calibrating the NR response.
@@
 An map normalizes the S1 signals to the center of the TPC volume, with correction factors averaging 8% in absolute value.
-S2 signals are normalized to the radial center of the liquid surface in the plane, in addition to being corrected for the depth-dependent probability of electron attachment to electronegative impurities in the LXe; the average values of these corrections are approximately 13% and 7%, respectively.
+S2 signals are normalized to the radial center of the liquid surface in the plane, in addition to being corrected for the measured electron lifetime; the average values of these corrections are approximately 13% and 7%, respectively.
 The responses are also normalized in time, using the tritium calibration period as the reference.
@@
 The DD data define an analogous “NR band,” for which nest is separately tuned to a similar level of precision via small adjustments in the NR yield parameters.
+For both the ER and NR models, the tuned parameters and the resulting light and charge yields are within the uncertainty bands reported in Ref.
+[21].
 The DD neutron interactions are confined to the top of the detector, where the average electric field is 7% higher than the center.
@@
 The other seven events were revealed as salt following the final definition of the data selection criteria and the likelihood models.
+More details about the salting distribution can be found in the Supplemental Material [24].
 The 1220 events which remain in the 3.3 tonne-year exposure of WS2024 after the application of all data selections and salt removal are shown in Figs.
@@
 Nuclear recoils that originate from spontaneous fission and reactions in detector materials can be vetoed by detecting the scattered neutrons and/or associated rays in the Skin and/or OD.
-Events vetoed by these systems are used as a sideband that is used to derive an in-situ constraint on the number of neutron events in the WIMP search data.
-As discussed later, the sideband is incorporated directly into the final statistical analysis; however, a sideband-only fit returns an expectation of SS counts, as shown in Table 1.
+Events vetoed by these systems are used as a sideband to derive an in-situ constraint on the number of neutron events in the WIMP search data.
+As discussed later, the sideband is incorporated directly into the final statistical analysis; a sideband-only fit returns an expectation of SS counts, as shown in Table 1.
 Secondary estimates from simulations based on detector material radioassays and from data using candidate MS neutron events tagged by the veto detectors return expectations of and neutron counts, respectively, consistent with the primary analysis.
@@
 Preliminary measurements of this effect for ECs in LZ were performed with sidebands of Skin-tagged and multiple scatter decays of 127Xe and 125Xe, isotopes that have been produced in situ via neutron activation.
-These yield a charge-yield ratio of for 5.2 keV L-shell EC decays relative to decays of the same energy.
+These yield a charge-yield ratio of for 5.2 keV L-shell EC decays relative to decays of the same energy [27].
 Two-neutrino DEC decays of 124Xe are observed in LZ with a half-life of [28], consistent with that in Ref.
@@
 As an independent measurement of the LL-shell charge yield is not available, the Thomas-Imel (TI) box model [31, 21] is invoked to relate the ionization density to the amount of electron-ion recombination at a given applied electric field.
-The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
+The ratio of the charge yield for LL-shell DEC to decays of the same energy, , is treated as a free parameter in the statistical analysis, bounded on one side by the measured for WS2024 (0.87) and on the other side by the result from a preliminary study of increasing the ionization density in the TI box model (0.65).
 The largest ER background contribution in this exposure comes from -emitting isotopes in the LXe bulk.
@@
 These are used to estimate the contributions of 39Ar and 85Kr, respectively.
-Exposure in the high-mixing state includes residual tritium and 14C activity following calibrations, and therefore both their beta spectra are incorporated into the background model for this period.
+Exposure in the high-mixing state includes residual tritium and 14C activity following their injection for ER calibrations, and both their beta spectra are therefore incorporated into the background model for this period.
 Radon emanates from materials exposed to xenon, dispersing in the active volume.
 With a measured FV activity of Bq/kg, 214Pb decays from the 222Rn chain, in particular those to the ground state, form the dominant ER background.
+We attribute the higher activity in WS2024 with respect to WS2022 (Bg/kg [25]) to the lower drift field, as charged radon daughter progeny move more slowly out of the fiducial volume.
 A method to identify events with high likelihood of being a 214Pb decay — the radon tag — is developed using simulations of the movement of daughter atoms and ions built from observed flow vectors between 222Rn-218Po -decay pairs.
-A tag volume in which it is probable to find a 214Pb decay is defined for each identified 218Po alpha decay, co-moving with the projected flow streamline for 81 min or approximately 3 the 214Pb half-life.
-The radon tag is applied in the low-mixing state, where the flow can be modelled, and when data are available for a continuous 81 min period such that tag volumes can be constructed.
-For WS2024, the radon tag identifies 15% of the tag-active exposure as 214Pb-rich, containing on average % of the 214Pb, estimated directly by measurements of its excited-state decays.
-The effective 214Pb activity in the untagged remainder of the tag-active exposure is thus Bq/kg.
-Both the radon-tagged and untagged data sets are used in the final WIMP analysis; the former can be seen alongside all the WIMP search data in the bottom panel of Fig.
-3.
-The separation into 214Pb-rich and 214Pb-poor exposures offers the benefit of being able to better identify and constrain other background contributions, 124Xe in particular.
+A tag volume in which it is probable to find a 214Pb decay is defined for each identified 218Po alpha decay, co-moving with the projected flow streamline for 81 min or approximately three times the 214Pb half-life.
+The tag volume includes two flowpaths to account for neutral and ionized daughters.
+The radon tag can only be applied in the low-mixing state, where the flow can be modeled.
+Additional periods in the low-mixing state during which radon tagging is not possible, for example around gaps in data acquisition or after flow disturbances induced by power outages, are labeled as radon-tag-inactive.
+To summarize, we have four categories of WS2024 exposure: (1) high-mixing, where a radon tag is not possible; (2) low-mixing, but with the radon tag inactive due to interruptions; (3) low-mixing, active radon tag and radon-tagged, with a high probability to contain 214Pb decays; and (4) low-mixing, active radon tag and radon-untagged, with a lower probability to contain 214Pb decays.
+For WS2024, 15% of the low-mixing, tag-active exposure is identified as radon-tagged, containing on average % of the total 214Pb as estimated directly by measurements of its excited-state decays.
+The effective 214Pb activity in the untagged remainder of the tag-active exposure is thus Bq/kg. The radon-tagged data set can be seen in the bottom panel of Fig. 3.
+The separation into radon-tagged and untagged exposures offers the benefit of being able to better identify and constrain other background contributions, 124Xe in particular.
 Figure 3: Black points show the final set of events passing all selection cuts with the panels distinguishing all events in WS2024 (top) from those that are radon tagged (bottom).
-The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band.
-Dark and light, gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively.
+The radon-tagged data set contains 60% and 15% of the full FV activity of the 214Pb and dispersed backgrounds, respectively; thus, the comparison with all data brings to prominence the non-214Pb backgrounds like 124Xe that are not well-described by the ER band.
+Dark and light gray and purple shading show the regions containing 68% and 95% of the ER portion of the background model and a 40 GeV/ WIMP, respectively.
 Green and orange contours show the same quantiles for the best-fit 124Xe distribution and the accidental coincidence events, respectively.
@@
 8.3 0
-55.8
+55.9
 8.2 0
@@
 19.4
-3.9
-21.4
-3.6
+2.5
+20.4
+2.4
 […]
@@
 91
-1203
-42 0
-The data are divided into six mutually exclusive samples, which are simultaneously fit using an unbinned, extended maximum likelihood.
-Five of these apply to the new 220 live-day exposure and describe: (1) events acquired in the high-mixing state; (2) events acquired in the low-mixing state when the radon tag was inactive; events acquired in the low-mixing state that are (3) radon tagged, and (4) radon untagged; and (5) events tagged by the Skin or OD.
+1202
+410
+The data are divided into six mutually exclusive samples, which are simultaneously analyzed using an unbinned, extended maximum likelihood.
+Five of these apply to the new 220 live-day exposure and describe: the (1-4) WS2024 samples described previously
+and (5) events tagged by the Skin or OD.
 The nuisance parameters, which float within their uncertainties, are the expected number of events for each background source, listed in Table 1; the radon tagging efficiency; the neutron veto tagging efficiency; and the
@@
 The sixth sample in the likelihood represents WS2022, containing the data and models from the first, 60 live-day WIMP search result [8].
+The full WS2024 data set is shown in the top panel of Fig.
+3.
+All six samples can be seen in Fig.
+S1 of the Supplemental Material [24], where we also report their exposures.
 Figure 4: Comparison between data and the best-fit model for WS2024 corresponding to the values shown in Table 1 for a background plus a 40 GeV/ WIMP signal fit in a projection along the ER band median, normalized to the band width.
@@
 upper limit.
-The upper and lower panels select the S1 ranges 3–40 phd and 40–80 phd with values of 0.91 and 0.27, respectively.
+The upper and lower panels select the S1 ranges 3–40 phd and 40–80 phd with values of 0.55 and 0.59, respectively.
 Shaded bands depict the central interval containing 68% of the combined systematic and statistical uncertainties of the model.
 The best-fit number of WIMPs at all masses tested (between 9 GeV/ and 100 TeV/) for the combined WS2024+WS2022 analysis is zero.
-The goodness of fit of the background-only model is assessed in binned 1D projections of the data and model into S1c, log10(S2c), reconstructed energy, distance from the ER band median, as well as the 2D space of .
+The goodness of fit of the background-only model is assessed in binned 1D projections of the data and model into , , reconstructed energy, distance from the ER band median, as well as the 2D space of .
 All samples show excellent model-data agreement, passing a Holm-Bonferroni test [32] with a significance level of 0.05.
@@
 Figure 5 shows the 90% confidence level upper limit on the spin-independent WIMP-nucleon cross section as a function of mass following inference performed with a two-sided, unbinned profile likelihood ratio test statistic [33].
-The limit is power constrained at all WIMP masses to below the median expectation following the recommendation of Ref.
+The limit is power constrained at all WIMP masses to below the median expectation following the recommendation in Ref.
 [*[][;Hereweemployapowerconstraintwhichusestheexpectedsensitivityoftheexperimentassumingnosignalispresent.]baxter2021recommended].
@@
 [8], which exhibit an underfluctuation in this region.
-The limit for the WS2024 data alone can be found in the Supplemental Material, with an approximate fluctuation across the entire mass range, primarily attributed to an underfluctuation of the accidental coincidence background in the region of largest overlap with expected WIMP distributions.
-The minimum of the limit curve is cm2 at 36 GeV/ while that of the median expected limit is cm2 at 40 GeV/c2.
+The limit for the WS2024 data alone can be found in the Supplemental Material [24], with an approximate fluctuation across the entire mass range, primarily attributed to an underfluctuation of the accidental coincidence background in the region of largest overlap with expected WIMP distributions.
+The minimum of the limit curve is cm2 at 40 GeV/ while that of the median expected limit is cm2 at 40 GeV/c2.
 The best median observation potential is cm2 at 40 GeV/c2.
@@
 Appendix A Appendix - Spin Dependent Results
-[…]
 Figure 6: The 90% confidence-level upper limit (solid-black curve) using the mean of the nuclear structure functions from [40], allowing for like-to-like comparison with other results.
@@
 The PICO-60 result relies on WIMP-scattering on the spin of the unpaired proton of 19F with minimal uncertainty.
+[…]
 The best-fit number of WIMP events is zero at all tested masses, from 9 GeV/ to 100 TeV/.
@@
 López Paredes et al. [2018] B. López Paredes, H. Araújo, F. Froborg, N. Marangou, I. Olcina, T. Sumner, R. Taylor, A. Tomás, and A. Vacheret, Astropart. Phys. 102, 56 (2018).
+Szydagis et al. [2025]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., Front. Detect. Sci. Technol. 2 (2025).
 Temples et al. [2021]D. J. Temples, J. McLaughlin, J. Bargemann, D. Baxter, A. Cottle, C. Dahl, W. Lippincott, A. Monte, and J. Phelan, Phys. Rev. D 104, 112001 (2021).
-de Laeter et al. [2003a]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure and Applied Chemistry 75, 683 (2003a).
+de Laeter et al. [2003]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure and Applied Chemistry 75, 683 (2003).
 Thomas and Imel [1987]J. Thomas and D. Imel, Physical Review A 36, 614 (1987).
-Szydagis et al. [2022]M. Szydagis, J. Balajthy, G. Block, J. Brodsky, E. Brown, J. Cutter, S. Farrell, J. Huang, E. Kozlova, C. Liebenthal, et al., arXiv:2211.10726 (2022).
-de Laeter et al. [2003b]J. R. de Laeter, J. K. Böhlke, P. D. Bièvre, H. Hidaka, H. S. Peiser, K. J. R. Rosman, and P. D. P. Taylor, Pure Appl. Chem. 75, 683 (2003b).
 Figure S1 shows the WS2024 and WS2022 events passing all data selection criteria grouped in their respective sample in the likelihood. Table S1 lists the exposure for each sample in tonne-years.
@@
 Appendix D WS2024-Only Limit
-Figure S2 shows the SI WIMP-nucleon cross section upper limit obtained using the data and models pertaining to the WS2024, 3.3 tonne year exposure only.
+Figure S2 shows the SI WIMP-nucleon cross section upper limit obtained using the data and models pertaining to the WS2024, 3.3 tonne-year exposure only.
 Figure S2: Upper limits (90% C.L.) on the spin-independent WIMP-nucleon cross-section as a function of WIMP mass from WS2024 (220.0 live days) only are shown with a solid black line, with a power constraint applied.
@@
 The green and yellow bands are the and sensitivity bands.
+Appendix F Parameters of the Salting Model
+As discussed in the main text, the salt is drawn from a distribution consisting of the sum of a WIMP-like exponential and a flat spectrum, as shown in Fig.
+S4.
+The parameter sets the maximum energy of the flat distribution and is drawn from a Gaussian distribution with mean keV and sigma keV.
+The parameter is chosen to mimic the recoil spectrum induced by a WIMP with mass , which is in turn drawn from a Gaussian with mean = 80 GeV/c2 and width = 20 GeV/c2.
+The normalization is drawn from a uniform distribution that spans from zero up to a rate that would be consistent with the upper limits from Ref.
+[8].
+Figure S4: Distribution from which the salt events are drawn. The parameters are described in the text.
+The actual normalization, exponential, and flat spectrum parameters used are still hidden to the collaboration as unveiling them would harm bias mitigation of ongoing analyses.
+The allowed ranges were chosen to ensure that salt events can plausibly span the WIMP-search energy range, and also reach the highest recoil energies that might be used in a future higher-energy analysis, such as for effective field theories of dark matter.